A loan is amortized over 7 years, with monthly payments at a nominal rate of 9.5% compounded monthly. The first payment is $1000, paid one month from the date of the loan. Each succeeding monthly payment will be 3% lower than the prior one. What is the outstanding balance immediately after the 30th payment is made?
Answer #1

Loan=1000/(1+9.5%/12)*(1-(0.97/(1+9.5%/12))^84)/(1-(0.97/(1+9.5%/12)))=25320.87240

Balance after 30th payment=25320.87240*(1+9.5%/12)^30-1000/(1+9.5%/12)*(1-(0.97/(1+9.5%/12))^30)/(1-(0.97/(1+9.5%/12)))*(1+9.5%/12)^30=9242.28745